OpenStudy (luigi0210):

Integration:

OpenStudy (luigi0210):

\[\LARGE \int~\frac{dx}{\sqrt{x^2-6x+13}}~dx\]

OpenStudy (rational):

complete the square for the quadratic inside bottom radical, then make a trig substitution

OpenStudy (luigi0210):

So something like? \[\LARGE \int~\frac{dx}{\sqrt{(x-3)^2+4}}~\] Then sub \(u=x-3\) and go from there>

OpenStudy (rational):

Yep ! or make the substitution directly in single step : sub \(\large x-3 = 2\tan u\)

OpenStudy (luigi0210):

Wait, *-4, not 4.

OpenStudy (rational):

+4 is right : 9 + 4 = 13

OpenStudy (luigi0210):

Oh, right, sorry, did that too quickly. Btw, isn't \(\Large \int~\frac{dx}{\sqrt{1+x^2}}=sinh^{-1}x\) not tan?

OpenStudy (rational):

that works too^

OpenStudy (luigi0210):

Alright, thank you rational :)

OpenStudy (rational):

earlier substitution is for ppl who are familiar wid hyperbolics yet, yw :)

OpenStudy (rational):

*not